OpenStudy (falco276):

HELP!!!!!!!!!!!!! Kim earns $30 for babysitting on Friday nights. She makes an average of $1.25 in tips per hour. Write the function of Kim's earnings and solve for how much she would make after 3 hours.

OpenStudy (falco276):

@CGGURUMANJUNATH

OpenStudy (falco276):

Here are the choices: y = 1.25x + 30; $33.75 y = 30x + 1.25; $43.75 y = 30x + 1.25; $91.25 y = 1.25x + 30; $31.25

OpenStudy (falco276):

@dumbcow @agent0smith @amistre64 @AllTehMaffs

OpenStudy (falco276):

@shamil98

OpenStudy (shamil98):

f(x) = 30 + 1.25x f(3) = 30 + 1.25(3)

OpenStudy (falco276):

So, it's A? :/

OpenStudy (shamil98):

yes..

OpenStudy (anonymous):

So we know that she earns $30 as long as she babysits, but that her tips are dependent on how many hours she works. So the formula for that would be F(x)=30+1.25x, where x is how many hours she works. Which answer works with that?

OpenStudy (shamil98):

30 is a non-changing value since its for one night, but 1.25 per hour , so that is multiplied to x ( the number of hours)

OpenStudy (falco276):

Thanks so much! :D

OpenStudy (falco276):

Hey I have more questions! Hold on. :)

OpenStudy (falco276):

Given the equation −2x − 13 = 8x + 7, which order of operations completely solves for x? Add 2x, then subtract 8x, lastly subtract 7 Add 2x, then add 13, lastly divide by 10 Subtract 8x, then add 13, lastly divide by −10 Subtract 8x, then add 2, lastly add 13

OpenStudy (anonymous):

This is asking you how would you solve for x. I would do that first and then see which answer matches what you did.

OpenStudy (falco276):

o-okay.